# Enhancing Systematic Decompositional Natural Language Inference Using Informal Logic

### v2

@@ Appendix C Example TreeWise Outputs
{{"index": 1, "factuality": [5, 5], "relevance": [5, 5], "redundancy": [5, 5], "complete_inference": 3, "explanation": "The decomposition does not specify the role of ’unlimited resources’ in the competition."}}
{{"index": 2, "factuality": [5, 5, 5], "relevance": [5, 1, 5], "redundancy": [5, 5, 5], "complete_inference": 2, "explanation": "The second premise weakens the hypotheses, the first premise is irrelevant, and the last premise doesn’t address unlimited resources"}}
{{"index": 3, "factuality": [5, 5], "relevance": [5, 5], "redundancy": [5, 5], "complete_inference": 2, "explanation": "The decomposition does not specify the role of ’unlimited resources’ in the competition."}}
{{"index": 4, "factuality": [5, 5, 5], "relevance": [5, 2, 5], "redundancy": [5, 5, 5], "complete_inference": 2, "explanation": "The decomposition does not specify the role of ’unlimited resources’ in the competition, and premise 2 is not very relevant"}}
@@
Document-conditioned forward generation prompt (continued):
@@ 4 DIFFERENT POSSIBLE DECOMPOSITIONS, 2 OR 3 FACTS EACH, ONE JSON ITEM PER LINE:
What is the role of decomposers in a food chain? (A) They consume other organisms. (B) They break down dead organic matter. (C) They use the Sun’s energy to make food. (D) They convert inorganic matter into organic matter.
@@ 20 DIFFERENT POSSIBLE DECOMPOSITIONS, 2 OR 3 FACTS EACH, ONE JSON ITEM PER LINE::
(on followup) how could we make these better? regenerate the 20 decompositions so that they are higher-fidelity and are better explanations for the hypothesis.
@@
Fact-conditioned decomposition generation prompt:
@@
A balloon filled with water is placed in a freezer. Which property of the water will change as the water reaches its freezing point? (A) color, (B) mass, (C) state, (D) weight
Which of these is not an instinctive behavior? (A) a bird building a nest, (B) a turtle burying its eggs, (C) a bear hibernating in winter, (D) a horse pulling a plowRefer to caption
@@ Appendix F Baselines
       $S\leftarrow\text{RetrieveSupportFacts}(q_{i})$ // Generate $m$ candidate trees using ICL Prompt
       $T_{1},\dots,T_{t}\leftarrow\text{GenerateCandidateTrees}(q_{i},S,m,t)$ foreach $T_{j}$ in $T$ do
@@
             $T_{j}\leftarrow\text{PruneTree}(T_{j})$ // Check for ungrounded leaves not in $S$
             $U\leftarrow\text{FindUngroundedLeaves}(T_{j},S)$ foreach $l_{k}$ in $U$ do
@@
             $F\leftarrow\{q_{i}\}$ $D\leftarrow[]$ $N\leftarrow 0$ while $F\neq\emptyset$ and $N<10$ do
                   $N\leftarrow N+1$ // Retrieve and flatten support facts for sentences in $F$
                   $S\leftarrow\text{Set}(\text{Flatten}(\text{RetrieveSupportFacts}(f)\text{ for }f\text{ in }F))$ // Generate one line of tree decomposition using ICL prompt
                   $d_{N}\leftarrow\text{GenOneStep}(q_{i},S,D)$ // Append line to decompositions
                   $D\leftarrow D+[d_{N}]$ // Create tree from decompositions
                   $T_{N}\leftarrow\text{CreateTree}(D)$ // Prune disconnected nodes from the tree
                   $T_{N}\leftarrow\text{PruneTree}(T_{N})$ // Check for ungrounded leaves
                   $U\leftarrow\text{FindUngroundedLeaves}(T_{N},S)$ foreach $l_{i}$ in $U$ do
@@
                               $D\leftarrow D+[\text{``}l_{i}\Leftarrow f\text{''}]$ ” ] $U\leftarrow U\setminus l_{i}$

### v3

@@ Appendix C Example TreeWise Outputs
{{"index": 1, "factuality": [5, 5], "relevance": [5, 5], "redundancy": [5, 5], "complete_inference": 3, "explanation": "The decomposition does not specify the role of unlimited resources in the competition."}}
{{"index": 2, "factuality": [5, 5, 5], "relevance": [5, 1, 5], "redundancy": [5, 5, 5], "complete_inference": 2, "explanation": "The second premise weakens the hypotheses, the first premise is irrelevant, and the last premise doesnt address unlimited resources"}}
{{"index": 3, "factuality": [5, 5], "relevance": [5, 5], "redundancy": [5, 5], "complete_inference": 2, "explanation": "The decomposition does not specify the role of unlimited resources in the competition."}}
{{"index": 4, "factuality": [5, 5, 5], "relevance": [5, 2, 5], "redundancy": [5, 5, 5], "complete_inference": 2, "explanation": "The decomposition does not specify the role of unlimited resources in the competition, and premise 2 is not very relevant"}}
@@
RDTE judgment prompt (continued):
@@
In the following exercise, I would like you to tell me if a line of reasoning is reasonable or not.
I will give you some facts and a possible conclusion. Please tell me whether the conclusion reasonably follows from the facts I gave you. If the conclusion does reasonably follow from the facts, then please answer "yes". If the conclusion does not reasonably follow from the facts, then please answer "no".
Note that some of the facts may be false, but I am only interested whether the conclusion would reasonably follow IF those facts were true. In other words, imagine a world in which the given facts are true. Would it be reasonable to draw the conclusion from those facts, if they were true?
Here are some examples:
IF Vegetables are plants.
AND Cabbages are plants.
THEN Cabbages are vegetables.
Q: Does the rules conclusion reasonably follow from the facts in the condition, if they were true? A: no
IF a nail is made of metal
AND metals conduct electricity
THEN a nail conducts electricity.
Q: Does the rules conclusion reasonably follow from the facts in the condition, if they were true? A: yes
IF dogs are birds
AND birds can fly
THEN dogs can fly
Q: Does the rules conclusion reasonably follow from the facts in the condition, if they were true? A: yes
IF sound requires matter to travel
AND a vacuum has no matter in it
THEN sound will not travel in a vacuum.
Q: Does the rules conclusion reasonably follow from the facts in the condition, if they were true? A: yes
IF Erosion can cause a landslide.
AND Mud is deposited by a landslide.
THEN Erosion can cause mud to be deposited.
Q: Does the rules conclusion reasonably follow from the facts in the condition, if they were true? A: yes
IF An animal needs to breathe in order to live.
AND Living things need water to live.
THEN Animals need water to live.
Q: Does the rules conclusion reasonably follow from the facts in the condition, if they were true? A: yes
IF Frogs also have a larynx, or voice box, to make sounds.
AND Animals that have vocal cords can make sounds.
THEN Frogs are animals.
Q: Does the rules conclusion reasonably follow from the facts in the condition, if they were true? A: no
IF All humans breathe.
AND Stones breathe.
THEN All humans and stones breathe.
Q: Does the rules conclusion reasonably follow from the facts in the condition, if they were true? A: yes
IF If a planet is rocky, it can only have a thin atmosphere.
AND Small planets and rocky planets have very thin atmospheres.
THEN If a planet is small and rocky, it has a thin atmosphere.
Q: Does the rules conclusion reasonably follow from the facts in the condition, if they were true? A: yes
IF Damming a river can cause a lake to form.
AND Dams are made of concrete.
THEN Dams are concrete lakes.
Q: Does the rules conclusion reasonably follow from the facts in the condition, if they were true? A: no
Now your turn!
{Entailment}
Answer the following yes/no question with either "yes" or "no". Just give a single word answer. Do not give any explanation or justification.
Here are some examples:
Is it true that an ocean contains large bodies of water? yes
Is it true that lightning is similar to a volcano erupting? no
Is it true that a fox squirrel is a kind of animal? yes
Is it true that a rainbow is a kind of electromagnetic discharge? no
Is it true that the surface of the moon is made up of water? no
Is it true that the surface of the moon is made up of gases? no
Is it true that a bluebird is a kind of animal? yes
Is it true that the moon s surface is made up of oceans? no
Is it true that the opposite of negative impact is positive impact? yes
Is it true that building a new highway through the area has a negative impact on the ecosystem? yes
Now lets do some more! Remember, answer with just a single word, yes or no.
[…]
Figure 10: BaRDa prompts from Clark et al. (2023) for entailment (upper) and factuality (lower) judgments.
@@ 4 DIFFERENT POSSIBLE DECOMPOSITIONS, 2 OR 3 FACTS EACH, ONE JSON ITEM PER LINE:
What is the role of decomposers in a food chain? (A) They consume other organisms. (B) They break down dead organic matter. (C) They use the Suns energy to make food. (D) They convert inorganic matter into organic matter.
@@ 20 DIFFERENT POSSIBLE DECOMPOSITIONS, 2 OR 3 FACTS EACH, ONE JSON ITEM PER LINE::
textbf{(on followup)}
how could we make these better? regenerate the 20 decompositions so that they are higher-fidelity and are better explanations for the hypothesis.
@@
Document-conditioned entailment prompt:
@@
A balloon filled with water is placed in a freezer. Which property of the water will change as the water reaches its freezing point? (A) color, (B) mass, (C) state, (D) weight Which of these is not an instinctive behavior? (A) a bird building a nest, (B) a turtle burying its eggs, (C) a bear hibernating in winter, (D) a horse pulling a plowRefer to caption
@@ Appendix F Baselines
       $S\leftarrow\text{RetrieveSupportFacts}(q_{i})$
       // Generate $m$ candidate trees using ICL Prompt
       $T_{1},\dots,T_{t}\leftarrow\text{GenerateCandidateTrees}(q_{i},S,m,t)$
       foreach $T_{j}$ in $T$ do
@@
             $T_{j}\leftarrow\text{PruneTree}(T_{j})$
             // Check for ungrounded leaves not in $S$
             $U\leftarrow\text{FindUngroundedLeaves}(T_{j},S)$
             foreach $l_{k}$ in $U$ do
@@
             $F\leftarrow\{q_{i}\}$
             $D\leftarrow[]$
             $N\leftarrow 0$
             while $F\neq\emptyset$ and $N<10$ do
                   $N\leftarrow N+1$
                   // Retrieve and flatten support facts for sentences in $F$
                   $S\leftarrow\text{Set}(\text{Flatten}(\text{RetrieveSupportFacts}(f)\text{ for }f\text{ in }F))$
                   // Generate one line of tree decomposition using ICL prompt
                   $d_{N}\leftarrow\text{GenOneStep}(q_{i},S,D)$
                   // Append line to decompositions
                   $D\leftarrow D+[d_{N}]$
                   // Create tree from decompositions
                   $T_{N}\leftarrow\text{CreateTree}(D)$
                   // Prune disconnected nodes from the tree
                   $T_{N}\leftarrow\text{PruneTree}(T_{N})$
                   // Check for ungrounded leaves
                   $U\leftarrow\text{FindUngroundedLeaves}(T_{N},S)$
                   foreach $l_{i}$ in $U$ do
@@
                               $D\leftarrow D+[\text{``}l_{i}\Leftarrow f\text{''}]$ ” ]
                               $U\leftarrow U\setminus l_{i}$